find the value of m that makes the proportion m over 4 = m-6 over 12 a true statement.
Answer by stanbon(75887)

find the value of m that makes the proportion m over 4 = m-6 over 12 a true statement.
-----------
Proportion:
m/4 = (m-6)/12
Cross-multiply to get:
12m = 4m-24
8m = -24
m = -3
--------------------
Cheers,
Stan H.